A circular pool is surrounded by a brick walkway 3 m wide. Find the radius of the pool if the area of the walkway is 198 m^2.
Answer by mananth(12270)

let the radius of the pool be x
radius of pool + walkway = x+3
..
area of pool + walkway = pi*(x+3)^2
area of pool = pi*x^2
..
area of walkway = 198 m^2
pi(x+3)^2-pi*x^2= 198
/pi
(x+3)^2-x^2=198/pi
x^2+6x+9-x^2=198/pi
6x+9 = 63
-9 -9
6x=63-9
6x=54
x= 9m radius of the pool
...
m.ananth@hotmail.ca